OpenStudy (anonymous):

simplify: (x^-1 + y^-1)/(x^-1*y^-1)

OpenStudy (anonymous):

(1/x + 1/y) /(1/x)(1/y) =(y+x)/(xy) * ( xy) =y+x=x+y

OpenStudy (anonymous):

\[\frac {(x^-1 + y^-1)}{(x^-1*y^-1)} = 1 \]

Directrix (directrix):

@Rohangrr OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct Give a man a fish and you feed him for a day. Teach a man to fish and you feed him for a lifetime. ~Chinese Proverb